Subject: APO 3.5m users committee minutes, 9/10/12
From: Suzanne Hawley
Submitted: Mon, 10 Sep 2012 22:18:46 -0700 (PDT)
Message number: 1288
(previous: 1287,
next: 1289
up: Index)
APO 3.5-m Users Committee Phonecon, 9/10/12
Attending: Suzanne Hawley, Mark Klaene, Al Harper, John Stocke (for John
Bally), Scott Anderson, Russet McMillan, Remy Indebetouw, Jon Holtzman,
Rene Walterbos, David Bowen (for Michael Strauss), and Bruce Gillespie.
John Wilson and Fred Hearty also attended.
Absent: Sean Moran
**********************************
User feedback and comments from institutional representatives:
o Chicago (Al Harper) - Al had nothing new to report.
o Washington (Scott Anderson) - Scott had nothing new to report.
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Princeton (Dave Bowen) - Dave had nothing new to report.
o Colorado (John Stocke) - John had nothing new to report.
o Virginia (Remy Indebetouw) - Remy said that a UVa user is interested in GFP
status and plans (see below).
o Johns Hopkins (Sean Moran) - No report.
GFP status and plans: Bill Ketzeback said that the direct imaging and IFS
modes will be supported in TUI, but the Fabry-Perot functions will not. The F-P
modes can be controlled through command lines within TUI, however. Russell Owen
will write a GFP user widget, after which most of the instrument should be
usable by the general community in TUI. Suzanne added that the PI group at
Goddard run the FP functions remotely, so it is possible to use it without
being at APO. The interested potential user at UVa should explore further with
Bruce Woodgate how this is done.
**********************************
Discussion of telescope/instruments report:
Mark noted that we were able to get some observing in during July, but the
summer monsoon made August difficult and the rains have lingered unusually late
into September. During the shutdown, M2 and M3 were realuminized, and M1 was
washed. On instrument topics, during the shutdown APO staff opened instrument
dewars for the first time without the instrument developers attending -- see
the report below for details. The new echelle grating was installed and
adjusted to compensate for its different blaze angle from the old grating;
wavelengths are significantly shifted on the detector by up to half an order.
We are not seeing the expected throughput improvements in the echelle, and
continued investigation as to why this is is underway. Echelle users must take
new calibrations and adjust their data reduction procedures for observations
taken with the new grating.
John Wilson asked if the DIS blue camera leak had been present for some time
(see below). Mark replied that it had been a problem for some years.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
05/30/12 through 09/3/12
0) Overview
Observing continued through much of the summer due to a shorter than normal
engineering shutdown this year. The weather in July allowed some observations
prior to the shutdown, but many scheduled observations were weathered out in
August. Shutdown went generally as planned and we returned to science on
schedule, although observing was hampered by rain, clouds, and humidity.
The summer shutdown also included major work on the emergency generator and the
computer room UPS. A new air conditioner was also added for the computer room.
1) Telescope
The secondary and tertiary mirrors were realuminized. The primary mirror was
washed in situ. Minor altitude axis work was also accomplished. Post-shutdown,
on-sky testing yielded nominal collimation, pointing, and tracking performance
2) Instruments -- This consumed most of the shutdown and resulted in venturing
into uncharted territory. Observatory staff for the first time opened and
serviced the internal dewars on several instruments without the benefit of the
developers being present. While a couple of minor issues were encountered,
this was a major success. The cryotiger compressors were also moved and fed
with cold air which should improve their performance.
DIS blue camera was opened up and serviced. Prior to opening, water
condensation was noticed on the flat-field corrector lens. A new Cryotiger
cold head was installed and all O-rings serviced. A new ion pump and vacuum
gauge were also installed. A likely source of the recent vacuum leaks was
found in a small chip in the Schmidt optics that crossed the o-ring seal. This
was sealed and appears to have fixed the vacuum leak.
DIS red camera was not planned to be opened, but it was also found to have
water condensate on the flat field corrector lens. As a result, the vacuum
seals were all cleaned and re-greased as well. Upon reassembly the corrector
lens received a small chip in the corner. This was due to the lack of
compliance in the lens mount, and was corrected by adding a spring mount. The
insulating disc between the cold head and cold strap was also found to be
cracked and was replaced with indium foil. Unfortunately this caused an
increase in bias noise that manifested itself like dark current. On August 27
we opened up the dewar and reinstalled the insulation disc and added braycote
and the camera returned to nominal performance.
Both cameras are running at near nominal summertime temperatures.
Agile had no issues this period.
SPIcam is still running 5-7 Deg K warmer than we would expect with a good
vacuum and new coldhead. The camera is ready for use. We are continuing to
investigate the temperature issue.
A new cart was delivered for TripleSpec which improves the mounting process and
appears to fix the issue of the instrument coming off the kinematic mounts.
There are still some interferences, so the cart is not in operation yet. These
should be fixed later this month.
Echelle is operational. The new ECam motherboard was installed to fix the
shutter issue. The echelle grating was also replaced. The old BL67 grating
was removed and a new MR152 grating installed in its place. Due to a slight
change in blaze angle the wavelengths have shifted. Total throughput is still
being characterized but initial data with cal lamps indicate little if any
improvement compared to Newport data for the two gratings.
NIC-FPS is operating in shared-risk observing mode due to occasional image
corruption issues. The ICC crashed at the end of shutdown. A replacement
drive was installed and the instrument is back to its nominal state.
**********************************
New Observing Specialist:
Mark announced that Gabrelle Saurage has left APO to pursue a degree in
Philosophy. Alysha Shugart, who has previous observing experience at
UT-Austin, has been hired to replace her. Alysha is presently undergoing
training, and is likely to solo with the telescope soon. Please welcome her
and be patient as she learns our observing procedures.
**********************************
Institutional passwords:
Mark said that we've been getting an increasing number of requests from users
for the operational passwords. The members of the Users Committee are being
sent these passwords regularly, and they should ensure that their users receive
them. Suzanne added that we should also send them to the institutional
schedulers as well, and Mark agreed.
**********************************
TripleSpec darks:
Mark received a question from a CU user regarding taking "true" darks with
TripleSpec. Since the slit wheel is inoperable, this is no longer an easy
option. TripleSpec operates in a differential mode, so darks are actually part
of the normal data acquisition and reduction. Mike Skrutskie and John Wilson,
the instrument developers, have commented that true darks are usually not
necessary. John Stock said he will pass this message back to the CU user, and
will suggest that he discuss it further with Mike Skrutskie.
**********************************
Science input for BoG meeting:
Suzanne mentioned that she and Mark need the science highlights for the annual
report earlier than usual, in September. This year we want them in Powerpoint
format, including figures with appropriate short captions, and brief text in
bullet form. We do not want cut-and-pasted pages of pdf text from a journal
article. Two or three highlights per institution are desired; more will be
considered as space allows.
**********************************
Upcoming ATI proposal options:
Suzanne said that we are always looking for new instrumental capabilities for
the 3.5-meter telescope. Virginia has developed some ideas to discuss here,
and other institutions are welcome to participate.
Fred outlined UVa's recent thoughts on an ATI proposal. The next ATI proposal
deadline is ~50 days away, so time is short. UVa is interested in a fiber link
between the APOGEE instrument and the 3.5-m telescope. This would have
advantages for using the new high-resolution H-band spectrograph on a bigger
telescope in dark time, which is when it is not being used for SDSS-III and
eventually AS3. UVa is also working on other proposals, so they are looking
for another institution to host the fiber feed proposal. The fiber feed from
the APOGEE instrument could have multi-fiber and/or IFU capabilities, using
technology currently being developed for the MaNGA survey in AS3. Also, the
system could be used for high-precision velocity studies, as APOGEE has already
demonstrated RV precision of a few tens of meters per seconds.
Rene asked if it would be possible to use a multi-fiber setup at the 3.5-m
using separate plug plates for different target fields. Fred said yes, and
there could also be "standard" cartridges for IFU configurations. These
plugable cartridges could be used at a side port with a rotator and guider, but
it would probably be simpler and better to use the NA2 port.
Another ATI proposal idea that UVa is considering, possibly for next year, is
to build a custom NIR Fabry-Perot module for NIC-FPS, replacing the Queensgate
unit that has never worked properly.
Jon Holtzman pointed out that ATI proposals need to have a strong science case,
and Suzanne said that ideally there would be interest from users at the
institutions who could participate in putting together such a case. There
followed an extended discussion on the politics, science justification, and
practicality for other institutions partnering on these, and other, instrument
initiatives. Suzanne decided that there would be an action item (see below) to
Users Committee members to follow up on the fiber feed proposal which is the
most advanced at this stage. Before we take this much further, we need to
discuss the proposal with the AS3 Director, Mike Blanton. Rene said that we
should consider offering community access of the APOGEE feed to the 3.5-m as a
possibility, especially considering the outcome of the NSF Portfolio Review and
also in our current bid to attract new ARC partners.
**********************************
Update on new partners:
Suzanne and Rene reported on the upcoming changes in the ARC partnership.
Princeton has formally announced that it is leaving the consortium in June 2013
although they will continue to lease time until June 2014. Chicago has
indicated that they intend to withdraw in the 2014-15 time frame. This will
open up the equivalent of a 33% share in the telescope. We are eager to get
names of institutions and contacts that we can approach as potential new ARC
partners. One thing that came from the Portfolio Review is that our relatively
cheap operational mode is an asset, as is our remote observing system. Also
noted by the review is our strong mountain staff and reliable operations,
compared to other mid-sized telescopes. If we also continue to keep our
priority to develop or obtain new instrumental capabilities, all of these
strengths will help in the engagement of new institutional partners.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Send new ARC partner suggestions to Suzanne and Rene (before next
meeting).
Status: Suzanne has received a small number of suggestions, and is
still looking for more. Please send your suggestions of an institution and a
contact person to her and Rene.
===> ACTION: Science slides for BOG meeting should be sent to Suzanne and Mark
before the end of September.
[new actions from this meeting]:
===> ACTION: WITHIN THE NEXT WEEK!! Users committee members should discuss
APOGEE fiber feed to 3.5m proposal with their users and forward feedback, and
ideally name(s) of interested users who would be willing to participate in an
ATI proposal to Suzanne, Mark and Fred Hearty (see email addresses on CC line
of this message).
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 1 October
at 8:30 AM Pacific Time. The agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1288 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO